# On Iiro Honkala’s contributions to identifying codes

### v2

@@ Abstract
Olivier Hudry Télécom Paris, France
Ville Junnila [^1] The author was funded in part by the Academy of Finland grant 338797. University of Turku, Finland
Antoine Lobstein CNRS, Université Paris-Saclay, France
@@ 1 Introduction
Theorem 2 gathers a lower bound from [62] and an upper bound from [5] for $Id_{r}(G)$, both sharp (see also [17] and [38]).
Theorem 2.
@@ 2 Complexity for general graphs
The decision problem associated with the computation of $Id_{r}(G)$ is NP-complete for any fixed $r\geqslant 1$, as stated by Theorem 3.
Theorem 3.
@@
The proof of Theorem 3 for $r=1$ is based on a polynomial transformation from the NP-complete problem 3-SAT. Let $(U;C_{1},C_{2},...,C_{m})$ be any instance of 3-SAT, where $U$ denotes the set of Boolean variables and where $C_{1},C_{2},...,C_{m}$ are the clauses, each clause containing three literals. The transformation uses two kinds of components. The first one, associated with each Boolean variable $x_{j}$ ($1\leqslant j\leqslant|U|$), contains six vertices named $x_{j}$, $\overline{x_{j}}$, $a_{j}$, $b_{j}$, $c_{j}$ and $d_{j}$ and the six edges $\{a_{j},b_{j}\}$, $\{b_{j},x_{j}\}$, $\{b_{j},\overline{x_{j}}\}$, $\{c_{j},d_{j}\}$, $\{c_{j},x_{j}\}$, $\{c_{j},\overline{x_{j}}\}$. The second one, associated with each clause $C_{i}$ ($1\leqslant i\leqslant m$), contains two vertices named $\alpha_{i}$ and $\beta_{i}$ and the edge $\{\alpha_{i},\beta_{i}\}$. Extra edges reflect the composition of any clause: if $C_{i}$ contains the three literals $x_{j}$, $x_{h}$ and $x_{k}$, we add the edges $\{\alpha_{i},x_{j}\}$, $\{\alpha_{i},x_{h}\}$, $\{\alpha_{i},x_{k}\}$. Figure 2 illustrates this transformation by specifying a part of the graph thus obtained.
@@ 3.2 Lower bounds
For the radius $r>1$, the first efforts to improve the lower bounds of Theorems 6 and 7 were made in [31]. This approach was further refined in [29].
@@ 3.3 Upper bounds
  • By Theorem 8, we have $Id_{1}(\mathbb{F}^{2})\geqslant 3$, $Id_{1}(\mathbb{F}^{3})\geqslant 4$ and $Id_{1}(\mathbb{F}^{5})\geqslant 10$. Moreover, the lower bounds can be attained by the constructions given in [62].
  • In [8]*, the lower bounds $Id_{1}(\mathbb{F}^{4})\geqslant 6$ and $Id_{1}(\mathbb{F}^{7})\geqslant 31$ of Theorem 8 are respectively improved to $7$ and $32$. Moreover, $1$-IdCs attaining the improved lower bounds are also given.
  • In [8]*, a $1$-IdC in $\mathbb{F}^{6}$ with $19$ codewords is presented. By Theorem 8, it is known that $Id_{1}(\mathbb{F}^{6})\geqslant 18$ and later, in [31], it is shown, based on extensive computer searches, that $Id_{1}(\mathbb{F}^{6})=19$.
For $n\geqslant 8$, the smallest cardinalities of $1$-IdCs remain unknown: the best known lower bounds follow by Theorem 8 and the best constructions have been given in [9]. For an efficient method of constructing 1-IdCs, we require the following definition of $\mu$-fold coverings.
Definition 3.
@@
  • In [7]*, it has been shown that $Id_{2}(\mathbb{F}^{4})=6$; besides giving a $1$-IdC with 6 codewords, the authors also improve the lower bound of Theorem 7 from $5$ to $6$. In the same paper, the constructions with $6$ and $8$ codewords are presented to attain the lower bounds $Id_{2}(\mathbb{F}^{5})\geqslant 6$ and $Id_{2}(\mathbb{F}^{6})\geqslant 8$ of Theorem 7, respectively.
@@
  • In [39]*, a $3$-IdC in $\mathbb{F}^{5}$ with $10$ codewords is given. Furthermore, in [29], the lower bound $Id_{3}(\mathbb{F}^{5})\geqslant 9$ of Theorem 7 is improved to $10$ by a careful analysis of the proof of the theorem in that particular case.
@@
For larger radii, the best known lower bounds mostly follow by Theorem 7 (and its refinement in [29]), when $n$ is large enough. The best known constructions are mainly due to [9], although progress has also been made in [28] and [31]. In addition, a construction with the asymptotically smallest possible cardinality is presented in the following theorem from [52]*.
Theorem 10 ([52]*).
@@ 3.4 Conjectures in binary Hamming spaces
In what follows, we concentrate more closely on the conjecture (ii). We first present some results concerning the weaker formulation $Id_{1}(\mathbb{F}^{n+1})\leqslant 2Id_{1}(\mathbb{F}^{n})$ of the conjecture. In [6]* and [8]*, first attempts towards this goal have been presented.
Theorem 11 ([8]*).
@@ 3.5 Identifying sets of vertices
First efforts on identifying sets of vertices in $\mathbb{F}^{n}$ were already made in the seminal paper [62]. In what follows, we first focus on the case with the radius $r=1$. The following theorem states that a $(1,\leqslant\ell)$-IdC is always a $(2\ell-1)$-fold $1$-covering. The theorem was presented in the case $\ell=2$ in [48]* and later generalized to the case $\ell>2$ in [68]. Its proof is based on an insightful application of Lemma 1.
Theorem 15 ([48]*, [68]).
@@
Inequality (2) of the previous theorem follows from [19, Chapter 14]*. The converse of the previous theorem also holds for $\ell\geqslant 3$ as shown in [63] (based again on Lemma 1).
Theorem 16 ([63]).
@@
Notice that the requirement $\ell\geqslant 3$ is essential for Theorem 16. Indeed, in [77]*, it has been shown that there exists an infinite family of $3$-fold $1$-covering codes which are not (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-identifying. However, in the same paper, optimal constructions for (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdCs have been presented based on the so-called $\pi(\mbox{$\bf u$})$-construction already applied in Theorem 9.
Theorem 17 ([77]*).
@@
In [77]*, it has been proved that there exist (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdCs $C_{1}\subseteq\mathbb{F}^{5}$ and $C_{2}\subseteq\mathbb{F}^{7}$ with $16$ and $48$ codewords, respectively. Hence, the codes $C_{1}$ and $C_{2}$ attain the lower bound of Theorem 15 implying that $C_{1}$ and $C_{2}$ are also perfect $3$-fold $1$-coverings. Applying the previous theorem to the codes $C_{1}$ and $C_{2}$, we obtain two infinite families of (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdCs which are perfect $3$-fold $1$-coverings attaining the lower bound of Theorem 15. Hence, an optimal (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdC is obtained for lengths $n\geqslant 5$ for which a perfect $3$-fold $1$-covering also exists (see  [19, Theorem 14.2.4]*). More precisely, an optimal (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdC in $\mathbb{F}^{n}$ is obtained for the lengths $n=3\cdot 2^{k}-1$ and $n=2^{k+2}-1$, where $k$ is a positive integer.
@@
Next, we consider lower bounds on $(r,\leqslant\ell)$-IdCs when the radius $r\geqslant 2$. Firstly, a lower bound similar to the one of Theorem 7 has been provided in [29, Theorem 3.4]. It gives the best known lower bounds on $(r,\leqslant\ell)$-IdCs when $r\geqslant 2$ and $n$ is relatively small compared to $r$ and $\ell$. For larger lengths $n$, the best known lower bounds are given by the following theorem due to [29] and [31].
Theorem 20 ([29], [31]).
@@
In what follows, we continue with the constructions of $(r,\leqslant\ell)$-IdCs when the radius $r\geqslant 2$. As discussed earlier, the cardinalities $Id_{(r,\leq\ell)}(\mathbb{F}^{n})$ are rather well-known for $r=1$ and $\ell\geqslant 2$. Hence, the approach of the following theorem by [31], which generalizes Theorem 14, is quite effective for constructing $(r,\leqslant\ell)$-IdCs.
Theorem 21 ([31]).
@@
Another approach for constructing $(r,\leqslant\ell)$-IdCs is based on a generalization of the direct sum construction of Theorem 18. This approach is studied in the following theorem, which has been presented in [29].
Theorem 22 ([29]).
@@
Most of the best known $(r,\leqslant\ell)$-IdCs are due to Theorems 21 and 22. However, some incremental improvements have been made in [25]. In addition, an asymptotic result similar to Theorem 10 for $\ell=1$ has been proved in [57].
Theorem 23 ([57]).
@@ 4 Infinite grids
Figure 3: Partial representations of the four grids: (a) the square grid $G_{S}$; (b) the triangular grid $G_{T}$: black vertices are codewords (cf. Theorem 26); (c) the king grid $G_{K}$; (d) the hexagonal grid $G_{H}$ (with two possible representations: as a honeycomb or as a brick wall).
@@ 4.1 The square grid
The square grid, $G_{S}$, has vertex set $V_{S}=\mathbb{Z}\times\mathbb{Z}$ and edge set $E_{S}=\{\{u,v\}:u-v\in\{(1,0),(0,1)\}\}$. The exact minimum density is known for $r=1$, as stated by Theorem 24. The value 7/20 as an upper bound for $\partial^{Id}_{1}(G_{S})$ comes from [18]* and from [4] as a lower bound. Figure 5 displays a construction providing the upper bound.
Theorem 24 ([4], [18]*).
@@
For $r\geqslant 2$, Theorem 25 specifies bounds for $\partial^{Id}_{r}(G_{S})$. The lower bound for $r=2$ comes from [59]; the general lower bound for $r\geqslant 3$ comes from [10]*; all the upper bounds come from [51]*.
Theorem 25.
@@ 5.1 Extremal values of usual parameters in twin-free graphs
The path $P_{n}$ provides the value of $f_{\delta}(r,n)$ for any $r$ and any $n\geqslant 2r+1$: the minimum of $\delta$ for a connected and $r$-twin-free graph is equal to 1. Case (b) below relies on the fact that the complete graph minus a maximum matching is $1$-twin-free (cf. Theorem 31). On the other hand, the minimum degree cannot be too large with respect to the ratio $n/r$, otherwise there will be $r$-twins; Theorem 33 specifies bounds stated thanks to this ratio.
Theorem 33.
@@
Theorem 35.
@@
Figure 7: The case $n=15$, $r=3$ of Theorem 35, Case (b); this graph is $3$-twin-free and has diameter ${r+1=4}$.
@@
In any connected graph of order $n$ which is not the complete graph, $\alpha$ lies between $2$ and $n-1$ (the star). Theorem 1 above, from [1], contributes to Case (b) and to the lower bound of Case (c) in the following theorem.
Theorem 36.
@@
The star gives Case (a) in Theorem 37.
Theorem 37.
@@ 5.2 Terminal graphs
Observe that, for $r>1$, $P_{2r+1}$ (which is 1-twin-free, see Theorem 1) is $r$-terminal: indeed, removing any vertex from $P_{2r+1}$ leaves a graph with $r$-twins. For $r=1$, the case of $P_{3}$ is particular: removing the middle vertex yields two isolated vertices, constituting a $1$-twin-free graph. This leads us to address the following questions:
@@
There is no 1-terminal graph. Indeed, as observed above, for $n=3$, $P_{3}$ is not 1-terminal. For $n\geqslant 4$, consider a vertex $v$ such that $V\setminus\{v\}$ is a 1-IdC of $G$ (by Theorem 2, such a vertex does exist). Then we can show that the graph $G\setminus\{v\}$, which may be connected or not, is still 1-twin-free and, hence, $G$ is not 1-terminal.
@@
We now come back to general graphs. Then, for $r\geqslant 3$, $P_{2r+1}$ is not the only $r$-terminal graph. One possible construction in order to obtain another $r$-terminal graph is the following: take a cycle of length $2r$ and add a neighbour to each vertex of the cycle but one; Figure 8 illustrates this construction for $r=3$. Theorem 40 is more specific:
Theorem 40.
@@
Observe that the construction of Theorem 40 does not work for $r=2$; the problem remains open:
Open problem 1.

### v3

@@ Abstract
volume: 191issue: 3-4
On Iiro Honkala’s Contributions to Identifying Codes
@@ 1 Introduction
Theorem 1.3 gathers a lower bound from [62] and an upper bound from [5] for $Id_{r}(G)$, both sharp (see also [17] and [38]).
Theorem 1.3
@@ 2 Complexity for general graphs
The decision problem associated with the computation of $Id_{r}(G)$ is NP-complete for any fixed $r\geqslant 1$, as stated by Theorem 2.1.
Theorem 2.1
@@
The proof of Theorem 2.1 for $r=1$ is based on a polynomial transformation from the NP-complete problem 3-SAT. Let $(U;C_{1},C_{2},...,C_{m})$ be any instance of 3-SAT, where $U$ denotes the set of Boolean variables and where $C_{1},C_{2},...,C_{m}$ are the clauses, each clause containing three literals. The transformation uses two kinds of components. The first one, associated with each Boolean variable $x_{j}$ ($1\leqslant j\leqslant|U|$), contains six vertices named $x_{j}$, $\overline{x_{j}}$, $a_{j}$, $b_{j}$, $c_{j}$ and $d_{j}$ and the six edges $\{a_{j},b_{j}\}$, $\{b_{j},x_{j}\}$, $\{b_{j},\overline{x_{j}}\}$, $\{c_{j},d_{j}\}$, $\{c_{j},x_{j}\}$, $\{c_{j},\overline{x_{j}}\}$. The second one, associated with each clause $C_{i}$ ($1\leqslant i\leqslant m$), contains two vertices named $\alpha_{i}$ and $\beta_{i}$ and the edge $\{\alpha_{i},\beta_{i}\}$. Extra edges reflect the composition of any clause: if $C_{i}$ contains the three literals $x_{j}$, $x_{h}$ and $x_{k}$, we add the edges $\{\alpha_{i},x_{j}\}$, $\{\alpha_{i},x_{h}\}$, $\{\alpha_{i},x_{k}\}$. Figure 2 illustrates this transformation by specifying a part of the graph thus obtained.
@@ 3.2 Lower bounds
For the radius $r>1$, the first efforts to improve the lower bounds of Theorems 3.4 and 3.5 were made in [31]. This approach was further refined in [29].
@@ 3.3 Upper bounds
  • By Theorem 3.6, we have $Id_{1}(\mathbb{F}^{2})\geqslant 3$, $Id_{1}(\mathbb{F}^{3})\geqslant 4$ and $Id_{1}(\mathbb{F}^{5})\geqslant 10$. Moreover, the lower bounds can be attained by the constructions given in [62].
  • In [8]*, the lower bounds $Id_{1}(\mathbb{F}^{4})\geqslant 6$ and $Id_{1}(\mathbb{F}^{7})\geqslant 31$ of Theorem 3.6 are respectively improved to $7$ and $32$. Moreover, $1$-IdCs attaining the improved lower bounds are also given.
  • In [8]*, a $1$-IdC in $\mathbb{F}^{6}$ with $19$ codewords is presented. By Theorem 3.6, it is known that $Id_{1}(\mathbb{F}^{6})\geqslant 18$ and later, in [31], it is shown, based on extensive computer searches, that $Id_{1}(\mathbb{F}^{6})=19$.
For $n\geqslant 8$, the smallest cardinalities of $1$-IdCs remain unknown: the best known lower bounds follow by Theorem 3.6 and the best constructions have been given in [9]. For an efficient method of constructing 1-IdCs, we require the following definition of $\mu$-fold coverings.
Definition 3
@@
  • In [7]*, it has been shown that $Id_{2}(\mathbb{F}^{4})=6$; besides giving a $1$-IdC with 6 codewords, the authors also improve the lower bound of Theorem 3.5 from $5$ to $6$. In the same paper, the constructions with $6$ and $8$ codewords are presented to attain the lower bounds $Id_{2}(\mathbb{F}^{5})\geqslant 6$ and $Id_{2}(\mathbb{F}^{6})\geqslant 8$ of Theorem 3.5, respectively.
@@
  • In [39]*, a $3$-IdC in $\mathbb{F}^{5}$ with $10$ codewords is given. Furthermore, in [29], the lower bound $Id_{3}(\mathbb{F}^{5})\geqslant 9$ of Theorem 3.5 is improved to $10$ by a careful analysis of the proof of the theorem in that particular case.
@@
For larger radii, the best known lower bounds mostly follow by Theorem 3.5 (and its refinement in [29]), when $n$ is large enough. The best known constructions are mainly due to [9], although progress has also been made in [28] and [31]. In addition, a construction with the asymptotically smallest possible cardinality is presented in the following theorem from [52]*.
Theorem 3.8 ([52]*)
@@ 3.4 Conjectures in binary Hamming spaces
In what follows, we concentrate more closely on Conjecture (ii). We first present some results concerning the weaker formulation $Id_{1}(\mathbb{F}^{n+1})\leqslant 2Id_{1}(\mathbb{F}^{n})$ of the conjecture. In [6]* and [8]*, first attempts towards this goal have been presented.
Theorem 3.9 ([8]*)
@@ 3.5 Identifying sets of vertices
First efforts on identifying sets of vertices in $\mathbb{F}^{n}$ were already made in the seminal paper [62]. In what follows, we first focus on the case with the radius $r=1$. The following theorem states that a $(1,\leqslant\ell)$-IdC is always a $(2\ell-1)$-fold $1$-covering. The theorem was presented in the case $\ell=2$ in [48]* and later generalized to the case $\ell>2$ in [68]. Its proof is based on an insightful application of Lemma 3.1.
Theorem 3.13 ([48]*, [68])
@@
Inequality (2) of the previous theorem follows from [19, Chapter 14]*. The converse of the previous theorem also holds for $\ell\geqslant 3$ as shown in [63] (based again on Lemma 3.1).
Theorem 3.14 ([63])
@@
Notice that the requirement $\ell\geqslant 3$ is essential for Theorem 3.14. Indeed, in [77]*, it has been shown that there exists an infinite family of $3$-fold $1$-covering codes which are not (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-identifying. However, in the same paper, optimal constructions for (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdCs have been presented based on the so-called $\pi(\mbox{$\bf u$})$-construction already applied in Theorem 3.7.
Theorem 3.16 ([77]*)
@@
In [77]*, it has been proved that there exist (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdCs $C_{1}\subseteq\mathbb{F}^{5}$ and $C_{2}\subseteq\mathbb{F}^{7}$ with $16$ and $48$ codewords, respectively. Hence, the codes $C_{1}$ and $C_{2}$ attain the lower bound of Theorem 3.13 implying that $C_{1}$ and $C_{2}$ are also perfect $3$-fold $1$-coverings. Applying the previous theorem to the codes $C_{1}$ and $C_{2}$, we obtain two infinite families of (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdCs which are perfect $3$-fold $1$-coverings attaining the lower bound of Theorem 3.13. Hence, an optimal (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdC is obtained for lengths $n\geqslant 5$ for which a perfect $3$-fold $1$-covering also exists (see  [19, Theorem 14.2.4]*). More precisely, an optimal (1,2)(1$\leqslant$ 2)( 1 , ⩽ 2 )-IdC in $\mathbb{F}^{n}$ is obtained for the lengths $n=3\cdot 2^{k}-1$ and $n=2^{k+2}-1$, where $k$ is a positive integer.
@@
Next, we consider lower bounds on $(r,\leqslant\ell)$-IdCs when the radius $r\geqslant 2$. Firstly, a lower bound similar to the one of Theorem 3.5 has been provided in [29, Theorem 3.4]. It gives the best known lower bounds on $(r,\leqslant\ell)$-IdCs when $r\geqslant 2$ and $n$ is relatively small compared to $r$ and $\ell$. For larger lengths $n$, the best known lower bounds are given by the following theorem due to [29] and [31].
Theorem 3.19 ([29], [31])
@@
In what follows, we continue with the constructions of $(r,\leqslant\ell)$-IdCs when the radius $r\geqslant 2$. As discussed earlier, the cardinalities $Id_{(r,\leq\ell)}(\mathbb{F}^{n})$ are rather well-known for $r=1$ and $\ell\geqslant 2$. Hence, the approach of the following theorem by [31], which generalizes Theorem 3.12, is quite effective for constructing $(r,\leqslant\ell)$-IdCs.
Theorem 3.20 ([31])
@@
Another approach for constructing $(r,\leqslant\ell)$-IdCs is based on a generalization of the direct sum construction of Theorem 3.17. This approach is studied in the following theorem, which has been presented in [29].
Theorem 3.21 ([29])
@@
Most of the best known $(r,\leqslant\ell)$-IdCs are due to Theorems 3.20 and 3.21. However, some incremental improvements have been made in [25]. In addition, an asymptotic result similar to Theorem 3.8 for $\ell=1$ has been proved in [57].
Theorem 3.22 ([57])
@@ 4 Infinite grids
Figure 3: Partial representations of the four grids: (a) the square grid $G_{S}$; (b) the triangular grid $G_{T}$: black vertices are codewords (cf. Theorem 4.3); (c) the king grid $G_{K}$; (d) the hexagonal grid $G_{H}$ (with two possible representations: as a honeycomb or as a brick wall).
@@ 4.1 The square grid
The square grid, $G_{S}$, has vertex set $V_{S}=\mathbb{Z}\times\mathbb{Z}$ and edge set $E_{S}=\{\{u,v\}:u-v\in\{(1,0),(0,1)\}\}$. The exact minimum density is known for $r=1$, as stated by Theorem 4.1. The value 7/20 as an upper bound for $\partial^{Id}_{1}(G_{S})$ comes from [18]* and from [4] as a lower bound. Figure 5 displays a construction providing the upper bound.
Theorem 4.1 ([4], [18]*)
@@
For $r\geqslant 2$, Theorem 4.2 specifies bounds for $\partial^{Id}_{r}(G_{S})$. The lower bound for $r=2$ comes from [59]; the general lower bound for $r\geqslant 3$ comes from [10]*; all the upper bounds come from [51]*.
Theorem 4.2
@@ 5.1 Extremal values of usual parameters in twin-free graphs
The path $P_{n}$ provides the value of $f_{\delta}(r,n)$ for any $r$ and any $n\geqslant 2r+1$: the minimum of $\delta$ for a connected and $r$-twin-free graph is equal to 1. Case (b) below relies on the fact that the complete graph minus a maximum matching is $1$-twin-free (cf. Theorem 5.1). On the other hand, the minimum degree cannot be too large with respect to the ratio $n/r$, otherwise there will be $r$-twins; Theorem 5.3 specifies bounds stated thanks to this ratio.
Theorem 5.3
@@
Figure 7: The case $n=15$, $r=3$ of Theorem 5.5, Case (b); this graph is $3$-twin-free and has diameter ${r+1=4}$.
Theorem 5.5
@@
In any connected graph of order $n$ which is not the complete graph, $\alpha$ lies between $2$ and $n-1$ (the star). Theorem 1.2 above, from [1], contributes to Case (b) and to the lower bound of Case (c) in the following theorem.
Theorem 5.6
@@
The star gives Case (a) in Theorem 5.7.
Theorem 5.7
@@ 5.2 Terminal graphs
Observe that, for $r>1$, $P_{2r+1}$ (which is 1-twin-free, see Theorem 1.2) is $r$-terminal: indeed, removing any vertex from $P_{2r+1}$ leaves a graph with $r$-twins. For $r=1$, the case of $P_{3}$ is particular: removing the middle vertex yields two isolated vertices, constituting a $1$-twin-free graph. This leads us to address the following questions:
@@
There is no 1-terminal graph. Indeed, as observed above, for $n=3$, $P_{3}$ is not 1-terminal. For $n\geqslant 4$, consider a vertex $v$ such that $V\setminus\{v\}$ is a 1-IdC of $G$ (by Theorem 1.3, such a vertex does exist). Then we can show that the graph $G\setminus\{v\}$, which may be connected or not, is still 1-twin-free and, hence, $G$ is not 1-terminal.
@@
We now come back to general graphs. Then, for $r\geqslant 3$, $P_{2r+1}$ is not the only $r$-terminal graph. One possible construction in order to obtain another $r$-terminal graph is the following: take a cycle of length $2r$ and add a neighbour to each vertex of the cycle but one; Figure 8 illustrates this construction for $r=3$. Theorem 5.10 is more specific:
Theorem 5.10
@@
Observe that the construction of Theorem 5.10 does not work for $r=2$; the problem remains open:
Open problem 1